Keller Cosmetics maintains an operating profit margin of 7% and asset turnover ratio of 2.
a. What is its ROA? (Enter your answer as a whole percent.)
  ROA %
b.

If its debt-equity ratio is 1, its interest payments and taxes are each $9,000, and EBIT is $25,000, what is its ROE? (Do not round intermediate calculations. Enter your answer as a whole percent.)

  ROE %
Answer #1

a). ROA = Operating Profit Margin * Asset Turnover Ratio = 7% * 2 = 14%

b). Net Income = EBIT - Interest Payment - Taxes = $25,000 - $9,000 - $9,000 = $7,000

EBT = EBIT - Interest Payment = $25,000 - $9,000 = $16,000

Debt Burden = Net Income / EBT = $7,000 / $16,000 = 0.4375

Equity Multiplier = Assets / Equity = [1 + 1] / 1 = 2

ROE = Equity Multiplier * ROA * Debt Burden

= 2 * 14% * 0.4375 = 12.25%